The half-life of Carbon-14 is 5,730 years. What is the probability that a randomly selected Carbon-14 atom will not yet have decayed in 4,000 years’ time?
Answer #1

here as life of an atom follows exponential distribution with parameter \beta =mean life =half life/ln(2)=5730/ln(2)

=8266.643 years

hence   probability that a randomly selected Carbon-14 atom will not yet have decayed in 4,000 years’ time

=P(X>4000) =e-x/\beta =e-4000/8266.643=0.6164